OpenStudy (anonymous):

Given the linear equation x + y = 5, find the slope of its graph. A. –5 B. –1 C. 1 D. 5 @zeinas.

OpenStudy (anonymous):

Hello again!

OpenStudy (anonymous):

Hi :). So let's put this in slope intercept form. That is y=mx+b. To do that, we need to put the x on the other side. Subtract x from both sides. What do you get?

OpenStudy (anonymous):

Um so how do I do that I subtract 5?

OpenStudy (anonymous):

You do it like this: x+y=5 Now subtract x from both sides -x+x+y=-x+5 Negative x and positive x cancel out. So what are you left with?

OpenStudy (anonymous):

y=5?

OpenStudy (anonymous):

wait y+5m+b?

OpenStudy (anonymous):

I meant y=5m+b

OpenStudy (anonymous):

Don't forget about the -x on the right side. So y=-x+5. Got that part?

OpenStudy (anonymous):

Oh yeah ok then what?

OpenStudy (anonymous):

It is now in slope intercept form. y=mx+b, where m is the slope, and b is the y intercept. In our case, 5 is b, and m is -1(when you have -x, that is the same as -1x).

OpenStudy (anonymous):

Hang on give me a sec

OpenStudy (anonymous):

The trick is to divide your coefficients. The slope would be 1.

OpenStudy (anonymous):

No it's not. The slope is -1.

OpenStudy (anonymous):

How is it -1?

OpenStudy (anonymous):

Sure about that?

OpenStudy (anonymous):

oh ok I get it now thanks I dont get how it would be 1 butI'll go with -1

OpenStudy (anonymous):

You're right... it's -A/B, therefore it would be -1/1, which would be -1.

OpenStudy (anonymous):

Np @British_am_I . Do you need help with anything else?

OpenStudy (anonymous):

Yeah I do can I tag you again?

OpenStudy (anonymous):

Your amazing help!!!!